OpenStudy (inowalst):

A right circular cylinder has a radius of 8 and a height of \(\pi ^{2}\). If a cube has the same volume as the cylinder, what is the length of an edge of the cube?

OpenStudy (anonymous):

\[vol cyl.=\pi(r)^2h=\pi(8)^2(\pi^2)=64 \pi^3=L^3...so.... L=4\pi\]

OpenStudy (anonymous):

volume of cylinder = \[\Pi \times 8^{2} \times \Pi ^{2}\] volme of cube = \[l \times l \times l = l ^{3}\] therefore \[l ^{3} = 64\Pi ^{3}\] \[l = \sqrt[3]{64\Pi ^{3}} = 4\]

OpenStudy (anonymous):

*l = 4 pi

OpenStudy (inowalst):

@tkhunny

OpenStudy (tkhunny):

Is there still a question? Excepting typos, both solutions look fine.

OpenStudy (inowalst):

Yes. I just dont understand it, so I was wondering if you could help me..

OpenStudy (tkhunny):

Volume of Right Circular cone is what? Given radius and height measurements... Volume of Cube is what? Given edge measurement... Both examples, above, show this pretty clearly. The ONLY difficulty I see with this problem is that the height is given in terms of \(\pi\). Try not to let that bother you. Just go with it.

OpenStudy (inowalst):

Okay.